OpenStudy (anonymous):

The film was a long one, lasting 206 minutes. How many hours and minutes long was the film?

hero (hero):

206/60 = 3.43 (ignore remainder) 3*60 = 180 206 - 180 = 26 The movie was 3 hours and 26 minutes long

OpenStudy (anonymous):

From Mathematica: QuotientRemainder[ 206, 60 ] = { 3, 26} -> 3 hours, 26 minutes

hero (hero):

robtobey, what's up with the re-posts? There's no need to re-post afterward. Just say you agree with the answer and move on or just give a medal for a correct response.

OpenStudy (anonymous):

Well in this case I don't see what harm was done. The calculations were slightly different in their approach. As far as medals are concerned I received zero.